The last dividend of the common stock of Acme Co. was $2.50 per share. Suppose the...

  1. The last dividend of the common stock of Acme Co. was $2.50 per share. Suppose the expected growth rate for the next two periods is 18% and 5% indefinitely after that and that you require a return of 15% for assets like this?
    1. Draw a timeline for this stock over 3 years, label the cashflows.
    2. What are the dividends paid in each of the next 3 years?
    3. What is the expected price of the stock in year 2?
    4. According to the dividend discount model, what is the value of a share of Acme stock?
Answer #1

D1 = 2.50(1.18) = $2.95

D2 = 2.50(1.18)2 = $3.48

D3 = 2.50(1.18)2(1.05) = $3.66

Using DDM Model,

Stock Price in year 2 = 3.66/(0.15 - 0.05)

Stock price in Year 2 = $36.60

Value of Stock = 2.95(1.15) + (3.48 + 36.60)/(1.15)2

Stock Price = $32.87